# SoREX: Towards Self-Explainable Social Recommendation with Relevant Ego-Path Extraction

### v1

@@ 1. Introduction
Recommender systems have become a common choice for alleviating information overload after the prosperity of the Internet (Herlocker et al., 2004; Li et al., 2025). However, they always suffer from the sparse interactions between users and items. According to social correlation theories (Cialdini and Goldstein, 2004), it is believed that users’ preferences can be influenced by their social relationships. With the proliferation of online social platforms, social recommendation has been designed to improve user-item interaction modeling, addressing data sparsity and cold start problem with the aid of user-user social regularizations (Jamali and Ester, 2010; Yang et al., 2016).
@@
Recently, Graph Neural Networks (GNNs) (He et al., 2020; Zhou et al., 2025) have been widely adopted in recommender systems owing to their robust ability to model structured data. Notably, contemporary social recommendation algorithms (Wu et al., 2019; Yu et al., 2021b; Tao et al., 2022; Hu et al., 2023; Ma et al., 2024) have embraced GNNs to harness high-order social context and collaborative information simultaneously. Despite the significant accuracy enhancements brought about by GNN-based methods, they often fall short in providing meaningful explanations for their predictions.
@@
However, the explainability of predictions made by social recommender systems is of paramount importance for both users and service providers. For users, explainability serves as a foundation for fostering engagement and trust in the system (Zhang et al., 2014), enabling them to make more informed decisions (Sarwar et al., 2001). Online A/B testing on e-commerce platforms has demonstrated that providing explanations can improve click-through rates in real-world commercial settings (Zhang et al., 2014). For service providers, explainability is equally critical. High-profile incidents, such as YouTube’s loss of advertising revenue due to the opaque nature of its recommendation algorithms (Hern, 2017), and Amazon’s discontinuation of a biased recruitment system (Dastin, 2018), underscore the potential risks and consequences of lacking transparency. Therefore, enhancing the explainability of recommendation models not only improves system transparency but also aids in uncovering systematic patterns, thereby deepening our understanding of underlying network characteristics (Zhang et al., 2020).
@@
Current research on explainable recommendation systems often draws on user reviews (Chen et al., 2018; Wang et al., 2018b; Tao et al., 2019) and knowledge graphs (KGs) (Wang et al., 2019; Yu et al., 2013; Xian et al., 2019). Review-based methods focus on generating coherent and persuasive explanations for users, but they heavily depend on the availability and quality of textual reviews (Chen et al., 2018). In many scenarios, such reviews may be sparse, biased, or even absent, especially for newly added items or users, which greatly limits the generalizability of these methods. On the other hand, KG-based methods aim to improve transparency by leveraging semantic paths within well-constructed knowledge graphs (Yu et al., 2013). However, such graphs are often unavailable or incomplete in practice, particularly in social networks or user–item interaction graphs, which restricts the applicability of KG-based approaches.
@@
To overcome the limitations of existing explanation methods, we emphasize substructure mining (Ying et al., 2019; Wang et al., 2021)—a widely used graph-based interpretability technique that extracts the most contextually relevant and interpretable subgraphs. This approach helps uncover complex, factor-specific patterns essential for understanding the reasoning behind recommendation rankings. In this work, we leverage the structural properties of user–item interaction graphs and social networks to enhance the transparency of social recommender systems, enabling more general and broadly applicable explanations. However, most substructure mining methods are post hoc (Ying et al., 2019; Luo et al., 2020; Shan et al., 2021), generating explanations only after model training. Such approaches are prone to distribution shifts (Zhang et al., 2023; Zhao et al., 2023), which can compromise explanation fidelity. As a result, recent efforts have focused on developing self-explainable GNNs (Wang et al., 2021; Miao et al., 2022a; Fan et al., 2024; Dai and Wang, 2024) that generate explanatory subgraphs as part of the prediction process. Nevertheless, existing methods primarily target graph-level classification and often fail to capture the complex, node-pair-specific substructures required for recommendation tasks. Moreover, explanations in recommendation typically answer the question, “Why should this user choose this item?”—yet in ranking-based settings, more persuasive explanations may arise from comparative reasoning, i.e., “Why should this item be ranked higher than others?” While (Yang et al., 2022) attempts to address this question, it relies heavily on textual reviews and does not explicitly improve model transparency.
@@
For clarity, we denote users’ multi-hop neighborhoods as their ego-nets. Since GNNs can only perceive the ego-net of the target user, we propose to extract a subgraph of the ego-net relevant to each individual candidate item and factor, serving as a candidate-aware and factor-specific explanation. To enable complex substructure investigation and provide detailed comparison, we aim to extract dense explanatory subgraphs instead of sparse ones, as suggested by (Miao et al., 2022b, a). We achieve this by transforming the ego-net into a set of ego-paths, shared by both towers, representing all multi-hop paths on the joint graph of the social network and user-item interaction graph originating from the target user. Random walk sampling ensures memory efficiency during ego-path generation. Next, we compute similarities between the given candidate and all observed ego-paths in each tower. Transforming these similarities into ego-path sampling probabilities, we obtain a subset of ego-paths relevant to each candidate and factor. Finally, explanation re-aggregation aggregates information from the sampled ego-path subsets into the final user representations used for downstream prediction, emphasizing candidate-aware and factor-specific neighborhood information. Ego-paths, forming motifs through interweaving, enable complex substructure investigation. Additionally, each candidate is assigned factor-specific explanation graphs with different structures and similarity distributions, naturally forming comparative relationships and explanations among candidate items.
@@ 2.1. Social Recommendation
Recommender systems have been widely explored as an effective approach for modeling user–item interactions (He et al., 2020; Li et al., 2024a; Zhou et al., 2019). However, the sparsity of user-item interaction often limits their performance. To alleviate this issue, social recommendation is proposed to incorporate social networks and enrich user–item interaction modeling. Early studies (Jamali and Ester, 2010; Ma et al., 2011; Wu et al., 2021; Guo et al., 2016; Yang et al., 2016) usually use social network for embedding regularization.
Recent works introduce prospering GNNs into social recommendation due to their strong capability in modeling relational data. GraphRec (Fan et al., 2019) is the pioneer to design social recommender with attention-based GNN. To model the recursive influence diffusion process, Diffnet (Wu et al., 2019) and Diffnet++ (Wu et al., 2020) adopt a multi-layer influence propagation architecture to learn the evolution of user preference. MHCN (Yu et al., 2021b) proposes a multi-channel hypergraph convolutional network to explicitly leverage high-order user relations. DESIGN (Tao et al., 2022) leverages multiple GNN encoders to learn from different factors, fusing their knowledge via knowledge distillation. SocialLGN (Liao et al., 2022) conducts message propagation in both social network and interaction graph in each GNN layer for knowledge fusion. ESRF (Yu et al., 2022) devises a GCN-based deep adversarial social recommendation framework. Jiang et al. (Jiang et al., 2024) identify the low preference homophily among socially connected users, and propose SHaRe to address the issue via social graph rewiring. Recent works also integrate self-supervised learning (Yu et al., 2021a; Wu et al., 2022; Hu et al., 2023) and graph denoising (Yang et al., 2024; Li et al., 2024b; Li and Wang, 2024) into GNN-based social recommendation.
@@ 2.2. Explainable Recommendation
Most explainable recommendation solutions generate explanations based on item profiles, user reviews and knowledge graphs (KG). Profile-based and review-based methods aim to help users understand their decisions via retrieved or generated textual information. Du et al. (Du et al., 2023) proposed a taste cluster based self-explainable collaborative filter method, using item tags for cluster description. As for review-based methods, the retrieval-based methods aim to select review text that matches prediction via attention (Chen et al., 2018), reinforcement learning (Wang et al., 2018a) and other feasible means. A recent work (Yang et al., 2022) also raises the conception of comparative explanation, but it is also a review-based method. On the other hand, generation-based methods aim to identify important aspects of interacted items (Wang et al., 2018b; Tao et al., 2019) or directly apply text generation techniques (Li et al., 2017; Yang et al., 2021) to synthesize explanations. KG based methods aim to improve the transparency of recommenders. They mostly employ multi-hop path reasoning on KG , and then apply the optimal paths as explanations. The difference among them mainly lies in the reasoning strategy, such as meta-path template based (Yu et al., 2013), LSTM based (Wang et al., 2019) and reinforcement learning based (Xian et al., 2019; Zhao et al., 2020) methods.
Existing profile-based, review-based, and knowledge graph (KG)-based explanation methods rely heavily on external data, such as user reviews or structured semantic triples. However, such data is often difficult to obtain, incomplete, or unavailable, especially in social recommendation scenarios involving new users or items. This limits their applicability in practice. While some prior work (Ji and Shen, 2016) introduces socially-aware explanations, it still fundamentally depends on reviews. In contrast, SoREX enhances transparency by leveraging only the intrinsic structure of user–item and user–user interactions, without requiring any auxiliary information. Furthermore, most KG-based methods focus on semantic reasoning and are ill-suited for modeling social influence. Our approach fills this gap by introducing a graph-based, self-explainable framework for social recommendation that supports comparative explanations based on mined substructures, thereby improving both interpretability and applicability in real-world settings.
@@ 2.3. Explainable GNN
The lack of interpretability of GNNs promote the development of GNN explainers. Early GNN explainers are post-hoc, generating posterior substructure-based explanations for trained GNNs. For example, GNNExplainer (Ying et al., 2019) and PGExplainer (Luo et al., 2020) generate edge masks or node feature masks to find the significant subgraphs; RGExplainer (Shan et al., 2021) searches for relevant subgraphs via reinforcement learning; XGNN (Yuan et al., 2020) trains a graph generator to generate explanation graphs; K-FactExplainer (Huang et al., 2024) proposes a factorized explainer to reflect one-to-many relationships between labels and explanatory substructures. However, recent works (Zhao et al., 2023; Zhang et al., 2023) prove that post-hoc explainers suffer from distribution shift between explanations and predictions.
Therefore, efforts have been made to develop self-explainable GNNs. SE-GNN (Dai and Wang, 2021) is one of the pioneers, using neighbors with the same label as the input node as explanation. ProtGNN (Zhang et al., 2022) and PxGNN (Dai and Wang, 2024) identifies representative graph patterns via prototype learning. ConPI (Wang et al., 2021) and ILP-GNN (Zhu et al., 2023) model similarity between neighbor sets of given node pair and infer the existence of link between them, which is explained by similar neighbors or neighbor pairs. GIB (Yu et al., 2020), GSAT (Miao et al., 2022a) and LRI (Miao et al., 2022b) all refer to information bottleneck (IB) theory. GIB proposes to train an explanatory graph generator with IB, while GSAT and LRI inject learnable randomness into GNN and sample dense explanatory graphs without spurious correlations. Some recent works have proposed further extensions to IB based methods. For example, PGIB (Seo et al., 2023) combines prototype learning with IB theory. There is also a group of causal learning based self-explainable GNNs (Sui et al., 2022; Fan et al., 2024) proposing to disentangle causal subgraphs from biased graph as explanations; Besides, a recent work GraphChef (Müller et al., [n. d.]) proposes to integrate decision tree into the message passing framework of GNNs to achieve self-explainability.
@@ 3.3. Friend Recommendation
Friend recommendation is critical for shaping and facilitating online social networks (Sankar et al., 2021; Song et al., 2022), helping users to connect with each other. We adopt it as an auxiliary task in SoREX. We can formulate friend recommendation problem as follows:
@@ 4.2. Basic Two-Tower Framework
To independently model social and user-item interaction factors and further lay the foundation for factor-specific explanation, we design a basic two-tower framework, which is composed of a social influence aware social tower and an interaction tower to learn from $\mathcal{G}_{s}$ and $\mathcal{G}_{r}$ respectively. Each tower is assigned with independent ID embeddings and GNN encoder, predicting its own ranking score. We define the ID embedding matrix in social and interaction tower as $\mathbf{E}^{s}\in\mathbb{R}^{(m+n)\times d}$ and $\mathbf{E}^{r}\in\mathbb{R}^{(m+n)\times d}$ respectively, where $d$ is the embedding dimension. For interaction tower, we adopt $k_{1}$-layer LightGCN (He et al., 2020) to encode user/item representations for prediction:
@@
As for the social tower, we adopt a mutant of LightGCN as the GNN encoder. Empirical studies in  (Tao et al., 2022) suggest that the influence of socially active users’ purchase decisions on which of other users has no significant increase or decrease compared with inactive users. Thus, the neighborhood aggregation function in LightGCN is not suitable for user-item interaction modeling in social recommendation. As suggested in  (Jiang et al., 2024), socially connected users will have more influence on each other if they are of similar purchase behaviors. Therefore, we devise a social influence aware aggregation function and integrate it into original LightGCN architecture to encode $\mathcal{G}_{s}$. The derivation of encoded user embedding $\mathbf{h}^{s}$ is formulated as follows:
@@ 4.3. Ego-Path Based Explanation
Although existing self-explainable GNNs are able to investigate complex explanatory substructures, they fail to provide comparative explanations for ranking predictions. To address these challenges, we propose to sample relevant subgraphs of $\mathcal{G}_{ego}$ for each candidate item $v_{j}$ and each concerned factor as candidate-aware and factor-specific explanation graph. This process is equivalent to sample subsets of $\mathcal{W}_{ego}$. Considering that the size of $\mathcal{W}_{ego}$ grows exponentially with the increase of hop number $k$, it is not feasible to perform path ranking over the whole $\mathcal{W}_{ego}$. Inspired by  (Yin et al., 2022), we randomly and uniformly sample $n_{w}$ walks originated from target user $u_{i}$ with maximum length $k$. Then we transform the sampled paths into the form of ego-paths introduced in section 3. This procedure essentially selects an ego-path subset $\hat{\mathcal{W}}_{ego}$ from all available ego-paths in $\mathcal{W}_{ego}$ to extract an even smaller explanation graph from. Note that repetitive walks will not be removed from $\hat{\mathcal{W}}_{ego}$ in order to implicitly reflect structural features of $\mathcal{G}_{ego}$.
@@
Then, we sample ego-path subsets $\tilde{\mathcal{W}}_{ego}^{r}(j)$ and $\tilde{\mathcal{W}}_{ego}^{s}(j)$ based on Bernoulli distribution $\beta_{t}^{r}\sim\text{Bern}(p_{t}^{r}(v_{j}))$ and $\beta_{t}^{s}\sim\text{Bern}(p_{t}^{s}(v_{j}))$ in interaction tower and social tower respectively. We only keep ego-paths with $\beta_{t}^{*}=1$ in corresponding tower. To make sure the gradient w.r.t. $p_{t}^{r}(v_{j})$ and $p_{t}^{s}(v_{j})$ is computable, we apply gumble-softmax reparameterization trick (Jang et al., 2016). To this end, we have obtained two ego-path subsets relevant to $v_{j}$ from the view of interaction and social factor respectively. They will be regarded as the factor-specific explanation graphs for $v_{j}$’s ranking result.

### v2

@@ 1. Introduction
Recommender systems have become a common choice for alleviating information overload after the prosperity of the Internet (cfevaluating; medretrieval). However, they always suffer from the sparse interactions between users and items. According to social correlation theories (cialdini2004social), it is believed that users’ preferences can be influenced by their social relationships. With the proliferation of online social platforms, social recommendation has been designed to improve user-item interaction modeling, addressing data sparsity and cold start problem with the aid of user-user social regularizations (socialmf; trustmf).
@@
Recently, Graph Neural Networks (GNNs) (lightgcn; tricolore) have been widely adopted in recommender systems owing to their robust ability to model structured data. Notably, contemporary social recommendation algorithms (diffnet; mhcn; design; cgcl; tail) have embraced GNNs to harness high-order social context and collaborative information simultaneously. Despite the significant accuracy enhancements brought about by GNN-based methods, they often fall short in providing meaningful explanations for their predictions.
@@
However, the explainability of predictions made by social recommender systems is of paramount importance for both users and service providers. For users, explainability serves as a foundation for fostering engagement and trust in the system (zhang2014explicit), enabling them to make more informed decisions (sarwar2001item). Online A/B testing on e-commerce platforms has demonstrated that providing explanations can improve click-through rates in real-world commercial settings (zhang2014explicit). For service providers, explainability is equally critical. High-profile incidents, such as YouTube’s loss of advertising revenue due to the opaque nature of its recommendation algorithms (hern2017youtube), and Amazon’s discontinuation of a biased recruitment system (dastin2018amazon), underscore the potential risks and consequences of lacking transparency. Therefore, enhancing the explainability of recommendation models not only improves system transparency but also aids in uncovering systematic patterns, thereby deepening our understanding of underlying network characteristics (zhang2020explainable).
@@
Current research on explainable recommendation systems often draws on user reviews (narre; mter; fact) and knowledge graphs (KGs) (kprn; KG-metapath; KGR-rl). Review-based methods focus on generating coherent and persuasive explanations for users, but they heavily depend on the availability and quality of textual reviews (narre). In many scenarios, such reviews may be sparse, biased, or even absent, especially for newly added items or users, which greatly limits the generalizability of these methods. On the other hand, KG-based methods aim to improve transparency by leveraging semantic paths within well-constructed knowledge graphs (KG-metapath). However, such graphs are often unavailable or incomplete in practice, particularly in social networks or user–item interaction graphs, which restricts the applicability of KG-based approaches.
@@
To overcome the limitations of existing explanation methods, we emphasize substructure mining (gnnexplainer; conpi)—a widely used graph-based interpretability technique that extracts the most contextually relevant and interpretable subgraphs. This approach helps uncover complex, factor-specific patterns essential for understanding the reasoning behind recommendation rankings. In this work, we leverage the structural properties of user–item interaction graphs and social networks to enhance the transparency of social recommender systems, enabling more general and broadly applicable explanations. However, most substructure mining methods are post hoc (gnnexplainer; pgexplainer; rgexplainer), generating explanations only after model training. Such approaches are prone to distribution shifts (mixup-explainer; post-hoc-consistent), which can compromise explanation fidelity. As a result, recent efforts have focused on developing self-explainable GNNs (conpi; gsat; disc; pxgnn) that generate explanatory subgraphs as part of the prediction process. Nevertheless, existing methods primarily target graph-level classification and often fail to capture the complex, node-pair-specific substructures required for recommendation tasks. Moreover, explanations in recommendation typically answer the question, “Why should this user choose this item?”—yet in ranking-based settings, more persuasive explanations may arise from comparative reasoning, i.e., “Why should this item be ranked higher than others?” While (compare-explain) attempts to address this question, it relies heavily on textual reviews and does not explicitly improve model transparency.
@@
For clarity, we denote users’ multi-hop neighborhoods as their ego-nets. Since GNNs can only perceive the ego-net of the target user, we propose to extract a subgraph of the ego-net relevant to each individual candidate item and factor, serving as a candidate-aware and factor-specific explanation. To enable complex substructure investigation and provide detailed comparison, we aim to extract dense explanatory subgraphs instead of sparse ones, as suggested by (lri; gsat). We achieve this by transforming the ego-net into a set of ego-paths, shared by both towers, representing all multi-hop paths on the joint graph of the social network and user-item interaction graph originating from the target user. Random walk sampling ensures memory efficiency during ego-path generation. Next, we compute similarities between the given candidate and all observed ego-paths in each tower. Transforming these similarities into ego-path sampling probabilities, we obtain a subset of ego-paths relevant to each candidate and factor. Finally, explanation re-aggregation aggregates information from the sampled ego-path subsets into the final user representations used for downstream prediction, emphasizing candidate-aware and factor-specific neighborhood information. Ego-paths, forming motifs through interweaving, enable complex substructure investigation. Additionally, each candidate is assigned factor-specific explanation graphs with different structures and similarity distributions, naturally forming comparative relationships and explanations among candidate items.
@@ 2.1. Social Recommendation
Recommender systems have been widely explored as an effective approach for modeling user–item interactions (lightgcn; ada; topic). However, the sparsity of user-item interaction often limits their performance. To alleviate this issue, social recommendation is proposed to incorporate social networks and enrich user–item interaction modeling. Early studies (socialmf; sr; cnsr; trustsvd; trustmf) usually use social network for embedding regularization.
Recent works introduce prospering GNNs into social recommendation due to their strong capability in modeling relational data. GraphRec (graphrec) is the pioneer to design social recommender with attention-based GNN. To model the recursive influence diffusion process, Diffnet (diffnet) and Diffnet++ (diffnet++) adopt a multi-layer influence propagation architecture to learn the evolution of user preference. MHCN (mhcn) proposes a multi-channel hypergraph convolutional network to explicitly leverage high-order user relations. DESIGN (design) leverages multiple GNN encoders to learn from different factors, fusing their knowledge via knowledge distillation. SocialLGN (sociallgn) conducts message propagation in both social network and interaction graph in each GNN layer for knowledge fusion. ESRF (esrf) devises a GCN-based deep adversarial social recommendation framework. Jiang et al. (share) identify the low preference homophily among socially connected users, and propose SHaRe to address the issue via social graph rewiring. Recent works also integrate self-supervised learning (sept; derec; cgcl) and graph denoising (gbsr; recdiff; gdssl) into GNN-based social recommendation.
@@ 2.2. Explainable Recommendation
Most explainable recommendation solutions generate explanations based on item profiles, user reviews and knowledge graphs (KG). Profile-based and review-based methods aim to help users understand their decisions via retrieved or generated textual information. Du et al. (ecf) proposed a taste cluster based self-explainable collaborative filter method, using item tags for cluster description. As for review-based methods, the retrieval-based methods aim to select review text that matches prediction via attention (narre), reinforcement learning (rl-extract) and other feasible means. A recent work (compare-explain) also raises the conception of comparative explanation, but it is also a review-based method. On the other hand, generation-based methods aim to identify important aspects of interacted items (mter; fact) or directly apply text generation techniques (nrt; saer) to synthesize explanations. KG based methods aim to improve the transparency of recommenders. They mostly employ multi-hop path reasoning on KG , and then apply the optimal paths as explanations. The difference among them mainly lies in the reasoning strategy, such as meta-path template based (KG-metapath), LSTM based (kprn) and reinforcement learning based (KGR-rl; KG-rl-2) methods.
Existing profile-based, review-based, and knowledge graph (KG)-based explanation methods rely heavily on external data, such as user reviews or structured semantic triples. However, such data is often difficult to obtain, incomplete, or unavailable, especially in social recommendation scenarios involving new users or items. This limits their applicability in practice. While some prior work (sr-explain) introduces socially-aware explanations, it still fundamentally depends on reviews. In contrast, SoREX enhances transparency by leveraging only the intrinsic structure of user–item and user–user interactions, without requiring any auxiliary information. Furthermore, most KG-based methods focus on semantic reasoning and are ill-suited for modeling social influence. Our approach fills this gap by introducing a graph-based, self-explainable framework for social recommendation that supports comparative explanations based on mined substructures, thereby improving both interpretability and applicability in real-world settings.
@@ 2.3. Explainable GNN
The lack of interpretability of GNNs promote the development of GNN explainers. Early GNN explainers are post-hoc, generating posterior substructure-based explanations for trained GNNs. For example, GNNExplainer (gnnexplainer) and PGExplainer (pgexplainer) generate edge masks or node feature masks to find the significant subgraphs; RGExplainer (rgexplainer) searches for relevant subgraphs via reinforcement learning; XGNN (xgnn) trains a graph generator to generate explanation graphs; K-FactExplainer (kfactexplainer) proposes a factorized explainer to reflect one-to-many relationships between labels and explanatory substructures. However, recent works (post-hoc-consistent; mixup-explainer) prove that post-hoc explainers suffer from distribution shift between explanations and predictions.
Therefore, efforts have been made to develop self-explainable GNNs. SE-GNN (se-gnn) is one of the pioneers, using neighbors with the same label as the input node as explanation. ProtGNN (protgnn) and PxGNN (pxgnn) identifies representative graph patterns via prototype learning. ConPI (conpi) and ILP-GNN (ilp-gnn) model similarity between neighbor sets of given node pair and infer the existence of link between them, which is explained by similar neighbors or neighbor pairs. GIB (gib), GSAT (gsat) and LRI (lri) all refer to information bottleneck (IB) theory. GIB proposes to train an explanatory graph generator with IB, while GSAT and LRI inject learnable randomness into GNN and sample dense explanatory graphs without spurious correlations. Some recent works have proposed further extensions to IB based methods. For example, PGIB (protgib) combines prototype learning with IB theory. There is also a group of causal learning based self-explainable GNNs (cal; disc) proposing to disentangle causal subgraphs from biased graph as explanations; Besides, a recent work GraphChef (graphchef) proposes to integrate decision tree into the message passing framework of GNNs to achieve self-explainability.
@@ 3.3. Friend Recommendation
Friend recommendation is critical for shaping and facilitating online social networks (grafrank; ssnet), helping users to connect with each other. We adopt it as an auxiliary task in SoREX. We can formulate friend recommendation problem as follows:
@@ 4.2. Basic Two-Tower Framework
To independently model social and user-item interaction factors and further lay the foundation for factor-specific explanation, we design a basic two-tower framework, which is composed of a social influence aware social tower and an interaction tower to learn from $\mathcal{G}_{s}$ and $\mathcal{G}_{r}$ respectively. Each tower is assigned with independent ID embeddings and GNN encoder, predicting its own ranking score. We define the ID embedding matrix in social and interaction tower as $\mathbf{E}^{s}\in\mathbb{R}^{(m+n)\times d}$ and $\mathbf{E}^{r}\in\mathbb{R}^{(m+n)\times d}$ respectively, where $d$ is the embedding dimension. For interaction tower, we adopt $k_{1}$-layer LightGCN (lightgcn) to encode user/item representations for prediction:
@@
As for the social tower, we adopt a mutant of LightGCN as the GNN encoder. Empirical studies in  (design) suggest that the influence of socially active users’ purchase decisions on which of other users has no significant increase or decrease compared with inactive users. Thus, the neighborhood aggregation function in LightGCN is not suitable for user-item interaction modeling in social recommendation. As suggested in  (share), socially connected users will have more influence on each other if they are of similar purchase behaviors. Therefore, we devise a social influence aware aggregation function and integrate it into original LightGCN architecture to encode $\mathcal{G}_{s}$. The derivation of encoded user embedding $\mathbf{h}^{s}$ is formulated as follows:
@@ 4.3. Ego-Path Based Explanation
Although existing self-explainable GNNs are able to investigate complex explanatory substructures, they fail to provide comparative explanations for ranking predictions. To address these challenges, we propose to sample relevant subgraphs of $\mathcal{G}_{ego}$ for each candidate item $v_{j}$ and each concerned factor as candidate-aware and factor-specific explanation graph. This process is equivalent to sample subsets of $\mathcal{W}_{ego}$. Considering that the size of $\mathcal{W}_{ego}$ grows exponentially with the increase of hop number $k$, it is not feasible to perform path ranking over the whole $\mathcal{W}_{ego}$. Inspired by  (subgraph_gnn), we randomly and uniformly sample $n_{w}$ walks originated from target user $u_{i}$ with maximum length $k$. Then we transform the sampled paths into the form of ego-paths introduced in section 3. This procedure essentially selects an ego-path subset $\hat{\mathcal{W}}_{ego}$ from all available ego-paths in $\mathcal{W}_{ego}$ to extract an even smaller explanation graph from. Note that repetitive walks will not be removed from $\hat{\mathcal{W}}_{ego}$ in order to implicitly reflect structural features of $\mathcal{G}_{ego}$.
@@
Then, we sample ego-path subsets $\tilde{\mathcal{W}}_{ego}^{r}(j)$ and $\tilde{\mathcal{W}}_{ego}^{s}(j)$ based on Bernoulli distribution $\beta_{t}^{r}\sim\text{Bern}(p_{t}^{r}(v_{j}))$ and $\beta_{t}^{s}\sim\text{Bern}(p_{t}^{s}(v_{j}))$ in interaction tower and social tower respectively. We only keep ego-paths with $\beta_{t}^{*}=1$ in corresponding tower. To make sure the gradient w.r.t. $p_{t}^{r}(v_{j})$ and $p_{t}^{s}(v_{j})$ is computable, we apply gumble-softmax reparameterization trick (gumble). To this end, we have obtained two ego-path subsets relevant to $v_{j}$ from the view of interaction and social factor respectively. They will be regarded as the factor-specific explanation graphs for $v_{j}$’s ranking result.